which is a proportion? 2/5 = 4/25, 4/6= 6/9, 4/3 = 6/8, 5/6 = 10/18
Answer by ikleyn(52778)

.
which is a proportion?
~~~~~~~~~~~~~~~~~~~~~~~~~

2/5 = 4/25    This is not

4/6 = 6/9     This is a proportion

4/3 = 6/8     This is not

5/6 = 10/18   This is not

On proportions read the lesson Proportions in this site, and also the lessons associated with it.